# Quadratic Chabauty Experiments on Genus 2 Bielliptic Modular Curves in the LMFDB

### v1

@@ Abstract.
We present results of quadratic Chabauty experiments on genus 2 bielliptic modular curves of Jacobian rank 2 that have recently been added to the LMFDB. We apply quadratic Chabauty methods over both the rationals and quadratic imaginary fields. In a number of cases, the experiments yielded algebraic irrational points among the set of mock rational points. We highlight specific notable examples, including the non-split Cartan modular curve $X_{ns}^{+}(15)$. Lastly, we offer a conjecture relating the level of the modular curve to the potential number fields over which points can arise.
@@ 1.1. Background and Motivation
This work has led to a wave of development already in the first nonabelian case, $n=2$, known as quadratic Chabauty. The quadratic Chabauty set $X(\mathbb{Q}_{p})_{2}$ is known to be finite when the rank of the Jacobian of the curve is less than $g-1+rk(NS(J))$, where $NS(J)$ denotes the Néron-Severi group of the Jacobian over $\mathbb{Q}$. Balakrishnan and Dogra establish this bound for finiteness in [BD18] and together with Müller compute some explicit examples. Developments have led to computational tools for new and more complicated examples in recent years. In [Bia20], Bianchi further developed quadratic Chabauty by using the $p$-adic sigma function instead of the double Coleman integral to compute the quadratic Chabauty locus for genus 2 bielliptic curves. Balakrishnan, Besser, Bianchi, and Müller later generalized this to rational points on genus 2 bielliptic curves over number fields [BBBM21]. Bianchi and Padurariu then implemented and scaled these methods to compute the set of rational points on bielliptic curves of Jacobian rank 2 from the LMFDB’s database of genus 2 curves [BP24, Cola].
@@ 1.2. Aim of this paper
    $H$ contains $-I$ and has genus at most 24 and level at most 335.
@@
    $H$ does not contain $-I$ and has genus at most 8 and level at most 335.
@@
In the rank 2 case, we vary the prime used for quadratic Chabauty to observe differences in what points appear. We take $p$ a prime of good, ordinary reduction for $X$ with $1003<p<$ 100. In the rank 1 case, we vary not only the prime but also the quadratic imaginary field used. We determine over which fields $\mathbb{Q}(\sqrt{-D})$ with $201<D<$ 20 and $D\in\{43,67,163\}$ the rank of the Jacobian of the curve is 2 (while we sought a tractable bound for $D$, adding the three larger values allows us to consider all imaginary quadratic fields of class number one at minimal additional cost). For each such field $\mathbb{Q}(\sqrt{-D})$, we restrict to primes $p$ of good, ordinary reduction that split in $\mathbb{Q}(\sqrt{-D})$.
@@
We note that this list of algebraic points is not provably exhaustive. We can, however, make conjectures based on the results of the explicit computations:
@@
Conjecture 1.5.
Let $X$ be a bielliptic modular curve of genus 2 and level $N$. If $Jac(X)$ is of rank 1 and $X$ has a mock rational point over $\mathbb{Q}(\sqrt{D})$ arise through quadratic Chabauty when computing over $\mathbb{Q}(\sqrt{D^{\prime}})$, then at least one of the following must hold:
@@
These hold for the genus 2 modular curves satisfying LMFDB Conditions (Section LMFDB Conditions). We provide code and data on GitHub to demonstrate computations in the paper [Fin25]. The code is modified from [Bal], which in turn is modified from [BPb] and uses SageMath [The24]. We will describe the new modifications in Sections 3 and 4.
@@ 1.3. Acknowledgements
Many thanks to Francesca Bianchi and Oana Padurariu for their virtual correspondence related to their previous work. We also thank Jennifer Balakrishnan for proposing the project and for helpful and insightful conversations throughout. The third modular curves workshop for the Simons Collaboration in Arithmetic Geometry, Number Theory, and Computation in March 2024, with support from the Simons Foundation, was also instrumental. Lastly, we are grateful to the reviewers of this article for their insight and attention to detail, both of which greatly improved the paper. The author was partially supported by NSF Grant DMS-1945452 and US-Israel BSF Grant 2022393.
@@ 2. Quadratic Chabauty for Bielliptic Curves
We have degree 2 maps $\phi_{i}:X\rightarrow E_{i}$ with $\phi_{1}(x,y)=(a_{6}x^{2},a_{6}y)$ and $\phi_{2}(x,y)=(a_{0}x^{-2},a_{0}yx^{-3})$.
@@
In particular, $W_{q}^{E}\subseteq\{0\}$ at all primes of good reduction for the given model of $E$.
@@
    $\Omega_{q}=(-W_{q}^{E_{1}}+W_{q}^{E_{2}}+\{-n\log q:-\text{ord}_{q}(a_{6})\leq n\leq\text{ord}_{q}(a_{0}),n\equiv 0\bmod 2\}$ )))
    $\cup(\log|a_{0}|_{q}-W_{q}^{E_{1}})\cup(-\log|a_{6}|_{q}+W_{q}^{E_{2}}$ )))
    and set $\Omega=\{\sum_{q\text{ bad}}w_{q}:w_{q}\in\Omega_{q}\}$. Then $\Omega$ is finite and contains $\widetilde{\rho}(X(\mathbb{Q}))$.
@@
In the development of the theory and the first examples of quadratic Chabauty for bielliptic curves, Balakrishnan and Dogra also define a set $\Omega$ [BD18, Theorem 1.2]. In this paper, the set of points whose image under a certain quadratic Chabauty function lands in $\Omega$ is a superset rather than strictly equal to $X(\mathbb{Q}_{p})_{2}$ (note that Balakrishnan and Dogra worked over general number fields). In turn, the set of Theorem 2.3 is a superset of the set of Balakrishnan and Dogra. As a result, the set $A$ of Equation 1.1 is a superset of $X(\mathbb{Q}_{p})_{2}$. This method of giving a superset that is explicitly computable is standard: in addition to [BD18], see [BDM+19], [BD21]. The authors of [BDM+19] remark that this can be thought of as an analogue to working with a suitable quotient of the Jacobian as in [Maz77] and [Mer96].
@@ 3.1. On Choice of Bielliptic Model
A given curve does not have a unique representation of the form in Equation 2.1. Most immediately, a change of variables $[x:y:z]\mapsto[-z:y:x]$ gives an isomorphism between $y^{2}=a_{6}x^{6}+a_{4}x^{4}+a_{2}x^{2}+a_{0}$ and $y^{2}=a_{0}x^{6}+a_{2}x^{4}+a_{4}x^{2}+a_{6}$. However, there are often other possibilities. For example, the following are isomorphic:
@@
$3ac^{5}a_{0}+2ac^{3}d^{2}a_{2}+acd^{4}a_{4}+bc^{4}da_{2}+2bc^{2}da_{4}+3bd^{5}a_{6}$ $=0$
$5a^{3}c^{3}a_{0}+a^{3}cda_{2}+3a^{2}bc^{2}da_{2}+a^{2}bd^{3}a_{4}+ab^{2}c^{3}a_{2}+3ab^{2}cd^{2}a_{4}+b^{3}c^{2}da_{4}+5b^{3}d^{3}a_{6}$ $=0$
$3a^{5}ca_{0}+a^{4}bda_{2}+2a^{3}b^{2}ca2+2a^{2}b^{3}da_{4}+ab^{4}ca_{4}+3b^{5}da_{6}$ $=0$
$e^{2}(c^{6}a_{0}+c^{4}d^{2}a_{2}+c^{2}d^{4}a_{4}+d^{6}a_{6})$ $\equiv 0\bmod(ad-bc)^{6}$
$e^{2}(15a^{2}c^{4}a_{0}+6a^{2}c^{2}d^{2}a_{2}+a^{2}d^{4}a_{4}+8abc^{3}da_{2}+8abcd^{3}a_{4}+b^{2}c^{4}a_{2}+6b^{2}c^{2}d^{2}a_{4}+15b^{2}d^{4}a_{6})$ $\equiv 0\bmod(ad-bc)^{6}$
$e^{2}(15a^{4}c^{2}a_{0}+a^{4}d^{2}a_{2}+8a^{3}bcda_{2}+6a^{2}b^{2}c^{2}a_{2}+6a^{2}b^{2}d^{2}a4+8ab^{3}cda_{4}+b^{4}c^{2}a_{4}+15b^{4}d^{2}a_{6})$ $\equiv 0\bmod(ad-bc)^{6}$
$e^{2}(a^{6}a_{0}+a^{4}b^{2}a_{2}+a^{2}b^{4}a_{4}+b^{6}a_{6})$ $\equiv 0\bmod(ad-bc)^{6}$
We can think of the first three as vanishing conditions: these ensure that the odd-power terms have zero coefficient. The latter three are integrality conditions that ensure the even-power terms are elements of $\mathbb{Z}$.
@@ 3.2. Execution of the Analysis
For example, consider the curve 16.48.2.a.1 with model $X:y^{2}=x^{6}-5x^{4}-5x^{2}+1$ of Jacobian rank 2. Working with both $p=11$ and $p=17$, we find $p$-adic points that appear to correspond to (±$(\pm 1,\pm 2\sqrt{-2})$. Without loss of generality, we let $Q=(1,2\sqrt{-2})$, since the other three points are redundant up to automorphism. One can immediately see that $Q$ is in $X(\mathbb{Q}(\sqrt{-2}))$ (as are its images under automorphism).
@@
Another change to the code relates to the computation of the set $\Omega$. When computing the local height at a given bad prime $q$, one can check if the curve has potential good reduction at $q$. If so, then the contribution at that prime can be understood entirely by information about the local height at a single $\mathbb{Q}_{q}$-point on the curve [BP24]. The original implementation assumes that such a point will exist. We remove this assumption by first checking in MAGMA [BCP97] if the curve is locally soluble at $q$ In the event that the code fails to obtain a $\mathbb{Q}_{q}$-point, we proceed to the general case to compute the height contribution at $q$.
@@ 3.3. Algebraic Weierstrass Mock Rational Points
    The prime $p$ satisfies the conditions for quadratic Chabauty for bielliptic curves: $p$ is of good and ordinary reduction for both $E_{i}$ where $Jac(X)\sim E_{1}\times E_{2}$, .
@@
Consider the curve $X$ with label 24.72.2.iy.1. It has model $y^{2}=6x^{6}+12x^{4}+12x^{2}+6$. The defining polynomial for $X$ factors over the rationals as $6(x^{2}+1)(x^{2}-x+1)(x^{2}+x+1)$, so we have three eligible terms in the factorization, and we expect to observe all 6 Weierstrass points. Note $E$ does not have ordinary reduction at 7 or 47, so we cannot use these primes for quadratic Chabauty. All other primes $p$ with $1003<p<$ 100 are of good, ordinary reduction. Since the Weierstrass points live over $\mathbb{Q}(i)$, and $\mathbb{Q}(\sqrt{-3})$, we can check the Legendre symbol of these primes with respect to $-1$ and $-3$. The behavior we see is as expected: the points (±$(\pm i,0)$ appear exactly when $p\equiv 1\bmod 4$ and the points (±$(\pm\frac{1}{2}\pm\frac{\sqrt{-3}}{2},0)$ appear exactly when $p\equiv 1\bmod 3$.
Next, consider the modular curve 16.48.2.a.1. We use $y^{2}=x^{6}-5x^{5}-5x^{2}+1$, which factors over the rationals as $(x^{2}+1)(x^{2}-2x-1)(x^{2}+2x-1)$. The first term has roots ±$\pm i$, and the latter terms have roots living over $\mathbb{Q}(\sqrt{2})$. The primes $p$ that are 1 modulo 4 are those such that $-1$ is a quadratic residue modulo $p$, so we consider the list $\{5,13,17,29,37,41,53,61,73,89,97\}$. We can check which primes are of good, ordinary reduction for the corresponding elliptic curves and obtain the list $\{17,41,73,89,97\}$. These are exactly the primes for which we observe the points (±$(\pm i,0)$. However, we do not observe the points over $\mathbb{Q}(\sqrt{2})$ for any prime $<100$. The conditions for a Weierstrass point to appear are necessary but not sufficient.
@@ 4. Computations over Quadratic Number Fields
    $\Omega_{q}=(-W_{q}^{E_{1}}+W_{q}^{E_{2}}+\{-n\log q:-\text{ord}_{q}(a_{6})\leq n\leq\text{ord}_{q}(a_{0}),n\equiv 0\bmod 2\}$ )))
    $\cup(\log|a_{0}|_{q}-W_{q}^{E_{1}})\cup(-\log|a_{6}|_{q}+W_{q}^{E_{2}}$ )))

### v2

@@ Abstract.
We present results of quadratic Chabauty experiments on genus 2 bielliptic modular curves of Jacobian rank 2 that have recently been added to the LMFDB. We apply quadratic Chabauty methods over both the rationals and quadratic imaginary fields. In a number of cases, the experiments yielded algebraic irrational points among the set of mock rational points. We highlight specific notable examples, including the non-split Cartan modular curve $X_{ns}^{+}(15)$. Lastly, we offer a conjecture relating the level of the modular curve to the potential number fields over which points can arise
@@ 1.1. Background and Motivation
This work has led to a wave of development already in the first nonabelian case, $n=2$, known as quadratic Chabauty. The quadratic Chabauty set $X(\mathbb{Q}_{p})_{2}$ is known to be finite when the rank of the Jacobian of the curve is less than g1+rk(NS(J))g-1+rk(NS(J)), where $NS(J)$ denotes the Néron-Severi group of the Jacobian over $\mathbb{Q}$. Balakrishnan and Dogra establish this bound for finiteness in [BD18] and together with Müller compute some explicit examples. Developments have led to computational tools for new and more complicated examples in recent years. In [Bia20], Bianchi further developed quadratic Chabauty by using the $p$-adic sigma function instead of the double Coleman integral to compute the quadratic Chabauty locus for genus 2 bielliptic curves. Balakrishnan, Besser, Bianchi, and Müller later generalized this to rational points on genus 2 bielliptic curves over number fields [BBBM21]. Bianchi and Padurariu then implemented and scaled these methods to compute the set of rational points on bielliptic curves of Jacobian rank 2 from the LMFDB’s database of genus 2 curves [BP24, Cola].
@@ 1.2. Aim of this paper
    $H$ contains I-I and has genus at most 24 and level at most 335.
@@
    $H$ does not contain I-I and has genus at most 8 and level at most 335.
@@
In the rank 2 case, we vary the prime used for quadratic Chabauty to observe differences in what points appear. We take $p$ a prime of good, ordinary reduction for $X$ with $3<p<100$. In the rank 1 case, we vary not only the prime but also the quadratic imaginary field used. We determine over which fields $\mathbb{Q}(\sqrt{-D})$ with $1<D<20$ and $D\in\{43,67,163\}$ the rank of the Jacobian of the curve is 2 (while we sought a tractable bound for $D$, adding the three larger values allows us to consider all imaginary quadratic fields of class number one at minimal additional cost). For each such field $\mathbb{Q}(\sqrt{-D})$, we restrict to primes $p$ of good, ordinary reduction that split in $\mathbb{Q}(\sqrt{-D})$.
We provide code and data on GitHub to demonstrate computations in the paper [Fin25]. The code is modified from [Bal], which in turn is modified from [BPb] and uses SageMath [The24]. We will describe the new modifications in Sections 3 and 4. From these computations, we obtain the following:
@@
We note that this list of algebraic points is not provably exhaustive. We can, however, make a conjecture based on the results of the explicit computations. It holds for the genus 2 modular curves satisfying LMFDB Conditions (Section LMFDB Conditions).
@@
We might hope for an analogous conjecture in the rank one case. Suppose $X$ is a bielliptic modular curve of genus 2 and level $N$ with $Jac(X)$ of rank 1. If $X$ has a mock rational point over $\mathbb{Q}(\sqrt{D})$ arise through quadratic Chabauty when computing over $\mathbb{Q}(\sqrt{D^{\prime}})$, we frequently observe one of the following:
@@
However, we have observed examples for which this is not the case. The curve 24.36.2.bo.1 has $\mathbb{Q}(\sqrt{-3})$-points appear when using $p=13$ and working over $\mathbb{Q}(\sqrt{-10})$. Also, the curve $X_{0}(37)$ has the aforementioned points over $\mathbb{Q}(\sqrt{-3})$ appear when using $p=13$ and working over $\mathbb{Q}(i)$. We suspect there is a systematic explanation for these not captured in the above divisibility conditions on $D$.
@@ 1.3. Acknowledgements
Many thanks are due to Francesca Bianchi and Oana Padurariu for their virtual correspondence related to their previous work. We also thank Jennifer Balakrishnan for proposing the project and for helpful and insightful conversations throughout. The third modular curves workshop for the Simons Collaboration in Arithmetic Geometry, Number Theory, and Computation in March 2024, with support from the Simons Foundation, was also instrumental. Lastly, we are grateful to the reviewers of this article for their insight and attention to detail, both of which greatly improved the paper. The author was partially supported by NSF Grant DMS-1945452 and US-Israel BSF Grant 2022393.
@@ 2. Quadratic Chabauty for Bielliptic Curves
We have degree 2 maps $\phi_{i}:X\rightarrow E_{i}$ with
(2.4) $\phi_{1}(x,y)=(a_{6}x^{2},a_{6}y),\hskip 14.22636pt\phi_{2}(x,y)=(a_{0}x^{-2},a_{0}yx^{-3}).$
@@
In particular, $W_{q}^{E}\subseteq\{0\}$ at all primes of good reduction for the given model of the elliptic curve $E$.
@@
    $\Omega_{q}=(-W_{q}^{E_{1}}+W_{q}^{E_{2}}+\{-n\log q:-\text{ord}_{q}(a_{6})\leq n\leq\text{ord}_{q}(a_{0}),n\equiv 0\bmod 2\}$ $)$
    $\cup(\log|a_{0}|_{q}-W_{q}^{E_{1}})\cup(-\log|a_{6}|_{q}+W_{q}^{E_{2}}$ $)$
[…]
@@
In the development of the theory and the first examples of quadratic Chabauty for bielliptic curves, Balakrishnan and Dogra also define a set $\Omega$ [BD18, Theorem 1.2]. In this paper, the set of points whose image under a certain quadratic Chabauty function lands in $\Omega$ is a superset rather than strictly equal to $X(\mathbb{Q}_{p})_{2}$ (note that Balakrishnan and Dogra worked over general number fields). In turn, the set of Theorem 2.3 is a superset of the set of Balakrishnan and Dogra, so the set $A$ of Equation 1.1 is a superset of $X(\mathbb{Q}_{p})_{2}$. This method of giving a superset that is explicitly computable is standard: in addition to [BD18], see [BDM+19], [BD21]. The authors of [BDM+19] remark that this can be thought of as an analogue to working with a suitable quotient of the Jacobian as in [Maz77] and [Mer96].
@@ 3.1. On Choice of Bielliptic Model
A given curve does not have a unique representation of the form in Equation 2.1. Most immediately, a change of variables $x:y:z]\mapsto[-z:y:x]$ gives an isomorphism between $y^{2}=a_{6}x^{6}+a_{4}x^{4}+a_{2}x^{2}+a_{0}$ and $y^{2}=a_{0}x^{6}+a_{2}x^{4}+a_{4}x^{2}+a_{6}$. However, there are often other possibilities. For example, the following are isomorphic:
@@
  • The following are all equal to zero:
    • $a_{0}(3ac^{5})+a_{2}(2ac^{3}d^{2}+bc^{4}d)+a_{4}(acd^{4}+2bc^{2}d)+a_{6}(3bd^{5})$.
    • $a_{0}(5a^{3}c^{3})+a_{2}(a^{3}cd+3a^{2}bc^{2}d+ab^{2}c^{3})+a_{4}(a^{2}bd^{3}+3ab^{2}cd^{2}+b^{3}c^{2})+a_{6}(5b^{3}d^{3})$.
    • $a_{0}(3a^{5}c)+a_{2}(a^{4}bd+2a^{3}b^{2}c)+a_{4}(2a^{2}b^{3}d+ab^{4}c)+a_{6}(3b^{5}d)$.
  • The following are all congruent to zero modulo $(ad-bc)^{6}$:
    • $e^{2}(c^{6}a_{0}+c^{4}d^{2}a_{2}+c^{2}d^{4}a_{4}+d^{6}a_{6})$
    • $e^{2}(15a^{2}c^{4}a_{0}+6a^{2}c^{2}d^{2}a_{2}+a^{2}d^{4}a_{4}+8abc^{3}da_{2}+8abcd^{3}a_{4}+b^{2}c^{4}a_{2}+6b^{2}c^{2}d^{2}a_{4}+15b^{2}d^{4}a_{6})$
    • $e^{2}(15a^{4}c^{2}a_{0}+a^{4}d^{2}a_{2}+8a^{3}bcda_{2}+6a^{2}b^{2}c^{2}a_{2}+6a^{2}b^{2}d^{2}a4+8ab^{3}cda_{4}+b^{4}c^{2}a_{4}+15b^{4}d^{2}a_{6})$
    • $e^{2}(a^{6}a_{0}+a^{4}b^{2}a_{2}+a^{2}b^{4}a_{4}+b^{6}a_{6})$
We can think of the first set as vanishing conditions: these ensure that the odd-power terms have zero coefficient. The latter are integrality conditions that ensure the even-power terms are elements of $\mathbb{Z}$.
@@ 3.2. Execution of the Analysis
For example, consider the curve 16.48.2.a.1, denoted $X$, with model given by $y^{2}=x^{6}-5x^{4}-5x^{2}+1$ of Jacobian rank 2. Working with both $p=11$ and $p=17$, we find $p$-adic points that appear to correspond to $(\pm 1,\pm 2\sqrt{-2})$. Without loss of generality, we let $Q=(1,2\sqrt{-2})$, since the other three points are redundant up to automorphism. One can immediately see that $Q$ is in $X(\mathbb{Q}(\sqrt{-2}))$ (as are its images under automorphism).
@@
Another change to the code relates to the computation of the set $\Omega$. When computing the local height at a given bad prime $q$, one can check if the curve has potential good reduction at $q$. If so, then the contribution at that prime can be understood entirely by information about the local height at a single $\mathbb{Q}_{q}$-point on the curve [BP24]. The original implementation assumes that such a point will exist. We remove this assumption by first checking in MAGMA [BCP97] if the curve is locally soluble at $q$ . In the event that the code fails to obtain a $\mathbb{Q}_{q}$-point, we proceed to the general case to compute the height contribution at $q$.
@@ 3.3. Algebraic Weierstrass Mock Rational Points
    The prime $p$ is of good and ordinary reduction for both elliptic curves $E_{i}$, where $Jac(X)\sim E_{1}\times E_{2}$, so it satisfies the conditions for quadratic Chabauty for bielliptic curves.
@@
Consider the curve $X$ with label 24.72.2.iy.1. It has a model given by $y^{2}=6x^{6}+12x^{4}+12x^{2}+6$. The defining polynomial for $X$ factors over the rationals as $6(x^{2}+1)(x^{2}-x+1)(x^{2}+x+1)$, so we have three eligible terms in the factorization, and we expect to observe all 6 Weierstrass points. Note $E$ does not have ordinary reduction at 7 or 47, so we cannot use these primes for quadratic Chabauty. All other primes $p$ with $3<p<100$ are of good, ordinary reduction. Since the Weierstrass points live over $\mathbb{Q}(i)$, and $\mathbb{Q}(\sqrt{-3})$, we can check the Legendre symbol of these primes with respect to 1-1 and 3-3. The behavior we see is as expected: the points $(\pm i,0)$ appear exactly when $p\equiv 1\bmod 4$ and the points $(\pm\frac{1}{2}\pm\frac{\sqrt{-3}}{2},0)$ appear exactly when $p\equiv 1\bmod 3$.
Next, consider the modular curve 16.48.2.a.1. We use $y^{2}=x^{6}-5x^{5}-5x^{2}+1$, which factors over the rationals as $(x^{2}+1)(x^{2}-2x-1)(x^{2}+2x-1)$. The first term has roots $\pm i$, and the latter terms have roots living over $\mathbb{Q}(\sqrt{2})$. The primes $p$ that are 1 modulo 4 are those such that 1-1 is a quadratic residue modulo $p$, so we consider the list $\{5,13,17,29,37,41,53,61,73,89,97\}$. We can check which primes are of good, ordinary reduction for the corresponding elliptic curves and obtain the list $\{17,41,73,89,97\}$. These are exactly the primes for which we observe the points $(\pm i,0)$. However, we do not observe the points over $\mathbb{Q}(\sqrt{2})$ for any prime <100<100. The conditions for a Weierstrass point to appear are necessary but not sufficient.
@@ 4. Computations over Quadratic Number Fields
    $\Omega_{q}=(-W_{q}^{E_{1}}+W_{q}^{E_{2}}+\{-n\log q:-\text{ord}_{q}(a_{6})\leq n\leq\text{ord}_{q}(a_{0}),n\equiv 0\bmod 2\}$ $)$
    $\cup(\log|a_{0}|_{q}-W_{q}^{E_{1}})\cup(-\log|a_{6}|_{q}+W_{q}^{E_{2}}$ $)$